The square of a positive number is seven more than six times the positive number. Find the number. Thans for helping.!!!!!!!!!!!!!!
Answer by checkley75(3666)

x^2=6x+7
x^2-6x-7=0
(x-7)(x+1)=0
x-7=0
x=7 answer
x+1=0
x=-1 answer
proof
7^2=6*7+7
49=42+7
49=49 and
-1^2=6*-1+7
1=-6+7
1=1